NEW HIGH RESOLUTION SPECTROSCOPY STUDIES OF METHYL NITRITE CH_{3}ONO

\underline{V. SIRONNEAU}, P. CHELIN, F. KWABIA TCHANA, I. KLEINER , Laboratoire Interuniversitaire des Systèmes Atmosphériques, CNRS et Universités Paris 7 et Paris 12, 61 av. Général de Gaulle, 94010, Créteil, France; J. ORPHAL, Institute for Meteorology and Climate Research, Karlsruhe Institute of Technology, 76344 Eggenstein-Leopoldshafen, Germany; O. PIRALI, Synchrotron SOLEIL, L’Orme des Merisiers Saint-Aubin, 91192 Gif-sur-Yvette, France; J-C. GUILLEMIN, Sciences Chimiques de Rennes -Ecole Nationale Suprieure de Chimie de Rennes -CNRS -35700 Rennes, France; L. MARGULES, R. MOTIYENKO, Laboratoire de Physique des Lasers, Atomes, et Molécules, UMR CNRS 8523, Universitée de Lille I, F-59655 Villeneuve dAscq Cédex, France; S. COOKE, W. J. YOUNGBLOOD, A. AGNEW, C. T. DEWBERRY, Department of Chemistry, University of North Texas, P.O. Box 305070, Denton, TX, 76203, U.S.A.

Methyl nitrite CH_{3}ONO is an important species in atmospheric chemistry involved in photochemical oxidation of volatile organic compounds. The cis conformer (more stable by about 298 cm^{−1}) has a high internal rotation potential barrier for the methyl group (731 cm^{−1})^{a} whereas for the trans conformer the barrier to internal rotation is extremely low (10 cm^{−1})^{b}, leading to large internal rotation splittings. Only one high resolution infrared study was performed prior to this study^{c}. For the first time, high-resolution spectrum of CH_{3}ONO was recorded in the far infrared region (30-500 cm^{−1}) using the synchrotron SOLEIL far-infrared beamline (AILES) and a Fourier transform (FT) spectrometer. Some 987 lines were assigned for the cis isomer up to J=65 and combined with 66 previously recorded microwave lines. In addition, high-resolution spectrum of the ν_{9} band of the cis isomer around 627.9 cm^{−1} was also recorded using the FT spectrometer at LISA. New microwave data is currently recorded to improve the knowledge of both the cis and trans ground state parameters.

aP. N. Gosh, A. Bauder and Hs. H. Gnthard, Chem. Phys. 53, 39-60 (1980)

bP. H. Turner, M. J. Corkill, and A. P. Cox, J. Chem. Phys. 83, 1473-1482 1979)

cL. M. Goss, C. D. Mortensen and T. A. Blake,J. Mol. Spectrosc., 225, 182-188 (2004)